# PatchTrack: A Comprehensive Analysis of ChatGPT’s Influence on Pull Request Outcomes

### v1

@@ 7.1 Threats to Validity (RQ1)
Another limitation stems from the representativeness of the dataset. Our analysis is restricted to pull requests where developers explicitly shared ChatGPT-generated content, meaning it does not capture private AI interactions or cases where ChatGPT was used without documentation in pull request discussions. This could introduce selection bias, as our findings are based only on visible ChatGPT contributions.
@@ 7.2 Trustworthiness of Qualitative Findings (RQ2, RQ3, RQ4)
To ensure credibility, we applied the Framework Method and thematic extraction via card sorting. Manual classification was performed in 277 cases ( PA, PN, NE, and CL), with Cohen’s Kappa of 0.85, indicating strong agreement among reviewers. Verbatim quotes from pull request discussions support data-driven interpretations, reducing researcher bias.
@@ 9 Conclusion
This study examined ChatGPT’s role in patch decision-making within pull requests by analyzing 285 merged pull requests and 53 closed pull requests from 255 repositories. We classified ChatGPT-generated patches as applied (PA), not applied (PN), or not suggested (NE), identifying 115 PA, 64 PN, and 106 NE cases. The median integration rate of 25% suggests that AI-generated patches are often treated as starting points rather than final implementations, and developers selectively modify or discard parts to fit project needs. Analysis of PN cases revealed six themes in which developers adapted AI suggestions to fit project constraints, while cases NE highlighted ChatGPT’s broader role in debugging, documentation, and conceptual guidance.
Our findings indicate that while ChatGPT accelerates software development by providing structured suggestions, human oversight remains critical in refining AI-generated contributions. Closed pull requests showed that rejections were often due to scope misalignment, maintainability concerns, or procedural barriers, underscoring the challenges of integrating AI into collaborative workflows. Overall, this study provides empirical insights into how generative AI is used in pull request workflows, informing developers, maintainers, and educators on best practices for AI-assisted coding and decision-making.
@@
Appendix A Supplementary Tables
| Classification Type | n = 1 | n = 2 | n = 3 | n = 4 |
| Patch Applied (PA) | 116 | 83 | 47 | 38 |
| Patch Not Applied (PN) | 63 | 96 | 132 | 141 |
| No Patch Exists (NE) | 106 | 106 | 106 | 106 |
| N-Gram | n = 1 | n = 2 | n = 3 | n = 4 | ||||||||||||
| Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | Acc. | Prec. | Recall | F1 | |
| PA | 97.2 | 96.6 | 95.6 | 96.1 | 95.6 | 95.3 | 88.4 | 91.7 | 95.9 | 88.1 | 88.1 | 88.1 | 96.7 | 90.9 | 85.7 | 88.2 |
| PN | 95.4 | 92.9 | 94.5 | 93.7 | 95.3 | 90.3 | 96.2 | 93.2 | 95.9 | 94.9 | 94.9 | 94.9 | 94.5 | 95.2 | 97.5 | 96.3 |
| NE | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
| All | 97.5 | 96.5 | 96.7 | 96.6 | 96.9 | 95.2 | 94.7 | 94.9 | 97.2 | 94.3 | 94.3 | 94.3 | 97.1 | 95.3 | 94.4 | 94.8 |

### v2

@@ 8.1 Threats to Validity (RQ1)
Another threat stems from the representativeness of the dataset. Our analysis is restricted to pull requests where developers explicitly shared ChatGPT-generated content, meaning it does not capture private AI interactions or cases where ChatGPT was used without documentation in pull request discussions. This could introduce selection bias, as our findings are based only on visible ChatGPT contributions.
A further threat involves the absence of a baseline comparison between ChatGPT-assisted and non-AI-assisted pull requests. Reviewers noted that it is difficult to assess the impact of AI without understanding how patch outcomes compare to traditional workflows. While this is a valid direction, comparing against non-AI PRs is complicated by the fact that many developers use AI tools without disclosing them. This makes it difficult to identify reliable control samples for comparison. One alternative is to use historical PRs from before the emergence of ChatGPT, but such comparisons may be confounded by temporal factors such as toolchain changes, developer norms, and review practices. We therefore focus on self-admitted ChatGPT usage (SACU), and leave broader comparisons as future work.
@@ 8.2 Trustworthiness of Qualitative Findings (RQ2, RQ3, RQ4)
To ensure credibility, we employed a rigorous qualitative analysis protocol based on the Framework Method and thematic extraction using card sorting [52, 114]. In RQ1, a sample of 230 pull requests (90 PA, 56 PN, and 84 NE) was manually validated to assess the accuracy of PatchTrack’s automated classification. This process, conducted independently by the two authors, followed a shared set of labeling criteria. Inter-rater agreement reached a Cohen’s Kappa of 0.85, indicating strong reliability.
For RQ1–RQ4, we applied a multi-stage qualitative process involving one author, a research assistant, and a senior researcher. Structured summaries were independently created for overlapping PR subsets and finalized through collaborative review. These summaries were then coded and consolidated into themes using card sorting. Discrepancies were resolved through consensus meetings involving the second author and the senior researcher. While we did not compute Cohen’s Kappa due to the interpretive nature of the analysis, agreement was established through iterative discussion and convergence on shared meanings. All findings were grounded in developer conversations and pull request excerpts, with examples and coding artifacts included in our replication package [42].
@@ 10 Conclusion
This study examined how developers engage with ChatGPT in pull request workflows by analyzing 338 real-world instances of self-admitted AI usage. By classifying cases into four categories (patch applied, patch not applied, no patch generated, and closed), we identified diverse ways developers adapt, reject, or conceptually reuse AI-generated suggestions.
Our analysis shows that full adoption of ChatGPT-generated patches is uncommon: the median integration rate was 25%. A qualitative analysis of applied patches revealed recurring patterns of structural integration, selective extraction, and iterative refinement, highlighting that developers often treat AI output as a starting point rather than a final solution. Even when no code is integrated, ChatGPT still influences pull requests through conceptual guidance, documentation revisions, and debugging strategies. In closed pull requests, we found both traditional barriers (e.g., code quality, scope misalignment) and AI-specific dynamics (e.g., discarded proofs-of-concept), underscoring that developer judgment remains central across all outcomes.
These findings highlight the collaborative nature of AI-assisted development and offer practical guidance for improving tool design, project workflows, and instructional practices. As generative AI continues to shape software engineering, understanding its nuanced role in human decision-making is essential for realizing its potential in real-world development contexts.